OpenStudy (anonymous):

Anyone in Algebra 1 Foundation? Connexus?

OpenStudy (anonymous):

Simplify each expression: 1.) 2d^28a^3= 2.) (5c-4)(-4m^2c^8)= 3.) (12x^2y^-2)^5(4xy^-3)^-7= 4.) Each square inch of your body has about 6.5*10^2 pores. Suppose the back of your head has an area of about 0.12*10^2 in^2. About how many pores are on the back of yoru head. Write your answer in Scientific Notation. Simplify each expreesion. 5.) 21x^3/3x^-1= 6.) 3c^3/e^5)^-4= 7.) 3.1x10^4/1.24x10^2= 8.) 5.1x10^5/1.7x10^2= Evaluate each function for the domain (1,2,3) 9.) f(x)=4^x= 10.) y=40(1/2)^x= Graph each function. 11.) f(x)=2.5^x= 12.)f(x)=1/2*3^x=

OpenStudy (abb0t):

That's too many questions. This is a site to help others, not simply give answers for medals. I will help yo uwith number one. You can start by factoring out a \(2\). \(\sf \color{red}{2}\)\((d^214a^3=0)\) You can then divide both sides by 2. To get final answer: \(d^214a^3\)

OpenStudy (anonymous):

Im in connexus algebra1foundations a . Can you message me and help me please? ;)